OpenStudy (anonymous):

HELP will fan and gold Find the slope of the line that passes through 9-2,0) and (5,-8)

OpenStudy (anonymous):

Is 9-2 correct or a typo?

OpenStudy (anonymous):

find the slope that passes through (-2,0) and (5,-8) sorry about the typo haha

OpenStudy (anonymous):

\[m = \frac{y2 - y1}{x2 - x1}\]\[m = \frac{-8 - 0}{5 - (-2)} = \frac{-8}{7}\]

OpenStudy (anonymous):

thank you!